Find the hypotenuse of each isosceles right triangle when the legs are of the given measure. Given=8

Answer by mananth(16949)

hypotenuse ^2 = leg1^2+leg2^2
H^2=8^2+8^2
H^2=64+64
H^2=128
H=sqrt 128
hypotenuse = 11.31 cm